A parallel plate capacitor with plate separation d is connected to a battery The capacitor is fully charged to O Coulombs and a voltage of V. (C is the capacitance and U is the stored energy.) Answer the following questions regarding the capacitor charged by a battery. For each statement below select True or False TrueAfter being disconnected from the battery, inserting a dielectric with K will decrease V FalseWith the capacitor connected to the battery, decreasing d increases U. True After being disconnected from the battery, decreasing d decreases TrueWith the capacitor connected to the battery, inserting a dielectric False After being disconnected from the battery, increasing d increases V C. with K will increase U FalseWith the capacitor connected to the battery, inserting a dielectric with k will increase C. Submit Answer Incorrect. Tries 1/99 Previous Tries
Answer #1

for part a) we have formula

C=Q/V=Q/Ed

so when we insert a dielectric ,electric field decreases because there is a capacitor

and this increases the capacitance and since capacitance is inversely proportional to V so V will decrease.

a) true

b) we have

C=Q/V=Q/Ed

with decrease in d capacitance increases

and U=1/2CV2

and C increases so energy increases

b) true

c) C=Q/Ed

with decrease in d , C increases as both of them are inversely proportional

c) false

d) again when we insert a dielectric the electric field decreases because there is a capacitor due to which the capacitance increases

as C=q/Ed

and we have U=1/2CV2

since u and c are directly proportional , so an increase in C will increase U

d) true

e) we have C=Q/Ed

with increase in d , C decreases

and C=Q/V

so decrease in C increases V

e)true

f) in presence of a capacitor the electric field decreases which increases capacitance

as C=Q/Ed

so answer

f) true.